OpenStudy (anonymous):

Find the volume of a sphere with a diameter of 4 feet. (Give your answer correct to 2 decimal places.)

OpenStudy (anonymous):

V = (4/3)*pi*(r^3) V = 33.51 cubic feet

OpenStudy (anonymous):

\[v=\frac{4 \pi r^3}{3} \]When r is 4, v = 268.08 The numerator starts off with 4^4, 256, and pi/3 is close to 1, so 33.51 is too small by a factor of 8tish.

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Sorry. theawesome's answer is correct. I used the diameter value 4, as the radius. The radius is 2.

OpenStudy (anonymous):

so the answer is 2

OpenStudy (anonymous):

The answer is:\[\frac{4 \pi 2^3}{3}=33.5103 \]